### v4

# Multiple Imputation for Nonresponse in Complex Surveys Using Design Weights and Auxiliary Margins

Abstract
Survey data typically have missing values due to unit and item nonresponse. Sometimes, survey organizations know the marginal distributions of certain categorical variables in the target population. As shown in previous work, survey organizations can leverage these distributions in multiple imputation for nonignorable unit nonresponse, generating imputations that result in plausible completed-data estimates for the variables with known margins. However, this prior work does not use the design weights for unit nonrespondents; rather, it relies on a set of fabricated weights for these units. We extend this previous work to utilize the design weights for all sampled units. We illustrate the approach using simulation studies.
Key Words: Item; Missing; Nonignorable; Unit
1 Introduction
Survey data usually suffer from both unit and item nonresponse. As a result, survey organizations have to make strong assumptions about the reasons for missingness, for example, the data are missing at random. One way to lessen reliance on assumptions is to utilize information in auxiliary data sources. For example, and pertinent to the setting of our work, survey organizations may know the population percentages or totals of some categorical variables in the survey. These could be available from recent censuses, administrative databases, or other high quality surveys (Sadinle and Reiter, 2017). Indeed, such information is frequently used in methods for handling survey nonresponse, such as calibration and raking.
In this short note, we consider settings where analysts seek to integrate auxiliary information on marginal distributions in multiple imputation for nonresponse (Rubin, 1987) in complex surveys. We build on the missing data with auxiliary margins, or MD-AM, framework introduced by Akande et al. (2021) and extended to complex surveys by Akande and Reiter (2022), Tang et al. (2024), and Yang and Reiter (2025). The latter three works impute missing values to ensure completed-data, survey-weighted inferences are plausible given the known margins. However, these works impose a simplifying condition on the unit nonrespondents, namely that the analyst does not use their design weights and instead replaces these weights with a convenient constant. This creates difficulties for design-based analyses, as the resulting weights for the full sample do not correspond to the design.
We propose MD-AM models that allow use of the design-based weights for all sampled units. The basic idea is as follows. We first impute all missing values due to item nonresponse using multiple imputation by chained equations (Van Buuren, 2018; Raghunathan et al., 2001). In each partially completed dataset, we then impute the unit nonrespondents’ values of the categorical variables that have auxiliary margins. To do so, we specify probabilities for imputations so that, in expectation, imputations result in plausible Horvitz and Thompson (1952) estimates in the completed data given the known margins, using the design weights for all sampled units. For the variables without auxiliary margins, we use the random hot deck imputation proposed in Yang and Reiter (2025).
The remainder of this note is organized as follows. Section 2 provides a brief review of the hybrid missing MD-AM model (Tang et al., 2024; Yang and Reiter, 2025), which we extend to incorporate design weights for unit nonrespondents. Section 3 introduces the methodology. Section 4 illustrates the methods using simulation studies. Codes are available at https://github.com/kevinxu47/MDAM.
2 Hybrid Missingness MD-AM Modeling
In reviewing the hybrid missingness MD-AM model, we closely follow the notation in Yang and Reiter (2025). Let $\mathcal{P}$ be a finite population comprising $N$ units. For $i=1,\dots,N$, let $\boldsymbol{x}_{i}=(x_{i1},\dots,x_{ip})$ denote the $p$ survey variables for unit $i$; let $I_{i}=1$ if unit $i$ is selected for inclusion in the survey and $I_{i}=0$ otherwise; let $\pi_{i}=\Pr(I_{i}=1)$ represent the probability that unit $i$ is selected into the survey; and, let $w_{i}^{d}=1/\pi_{i}$ be its design weight. Let $n=\sum_{i=1}^{N}I_{i}$ be the intended sample size of the survey. We refer to the sampled units as $\mathcal{S}$.
Let $U$ be the unit nonresponse indicator so that, for each unit $i\in\mathcal{S}$, $U_{i}=1$ if the unit provides no responses to any survey questions and $U_{i}=0$ otherwise. Let $\boldsymbol{R}=(R_{1},\dots,R_{p})$ be item nonresponse indicators corresponding to $\boldsymbol{X}=(X_{1},\dots,X_{p})$ so that, for any unit $i$ with $U_{i}=0$ and any variable $X_{j}$, $R_{ij}=1$ if unit $i$ does not respond to question $X_{j}$ and $R_{ij}=0$ otherwise. When $U_{i}=1$, $(R_{i1},\dots,R_{ij})$ are undefined.
The survey organization has auxiliary information $\mathcal{A}$ on the marginal distributions of a subset of the $p$ survey variables. For $j=1,\dots,p$, let $\mathcal{A}_{j}$ comprise the known marginal for $X_{j}$ when it exists. We write $X_{j}\in\mathcal{A}$ whenever $X_{j}$ has a known margin in $\mathcal{A}$. For convenience, we presume each $\mathcal{A}_{j}$ comprises population totals; incorporating percentages is a simple modification. For any categorical $X_{j}$ taking on levels $c=\{1,\dots,m_{j}\}$ and $X_{j}\in\mathcal{A}$, let $T_{jc}=\sum_{i=1}^{N}I(x_{ij}=c)$ be the total number of units in $\mathcal{P}$ at level $c$. Here, $I(\cdot)=1$ when the condition inside the parenthesis is true and $I(\cdot)=0$ otherwise.
The hybrid missingness MD-AM model specifies a joint distribution of $(\boldsymbol{X},\boldsymbol{R},U)$. It uses a pattern mixture model for unit nonresponse coupled with selection models for item nonresponse. Specifically, the model presumes $U\sim\textrm{Bernoulli}(\pi_{u})$, where $\pi_{u}=\Pr(U=1)$ is the marginal probability of unit nonresponse. For any $j=1,\dots,p$, let $g_{j}(-)$ represent the model for $X_{j}$ given $X_{1},\dots,X_{j-1}$, and let $\Omega_{j}$ and $\theta_{j}$ be the corresponding model parameters. For example, $g_{j}(-)$ could be a logistic regression of $X_{j}$ on some function of $X_{1},\dots,X_{j-1}$, possibly including interactions, as well as a main effect for $U$ when $X_{j}\in\mathcal{A}$. Thus, for $\boldsymbol{X}|U$, the hybrid missingness MD-AM model uses
$X_{1}\mid U$ $\sim$ $g_{1}(\Omega_{1},\theta_{1}UI(X_{1}\in\mathcal{A})$ (2.1)
$X_{j}\mid X_{1},\dots,X_{j-1},U$ $\sim$ $g_{j}(X_{1},\dots,X_{j-1},\Omega_{j},\theta_{j}UI(X_{j}\in\mathcal{A}),\textrm{ for }j=2,\dots,p.$ (2.2)
Because the distribution for any $X_{j}\in\mathcal{A}$ differs for unit nonrespondents and respondents when $\theta_{j}\neq 0$, the model encodes potentially missing not at random mechanisms for unit nonresponse. The key identifying assumption is that the predictor function for any $X_{j}\in\mathcal{A}$ not include interactions between $U$ and elements of $(X_{1},\dots,X_{j-1})$. This assumption is a version of the additive nonignorable missingness mechanism (Hirano et al., 2001; Sadinle and Reiter, 2019).
For the model for each $R_{j}$, let $h_{j}(-)$ be some predictor function that excludes the corresponding $X_{j}$ but may include main and interaction effects involving other variables in $\boldsymbol{X}$. For $j=1,\dots,p$, we have
$R_{j}\mid X_{1},\dots,X_{p},U=0\sim\text{Bernoulli}(\pi_{R_{j}}),\,\,\,\text{logit}(\pi_{R_{j}})=h_{j}(X_{1},\dots,X_{j-1},X_{j+1},\dots X_{p},\Phi_{j}).$ (2.3)
The models in (2.3) encode itemwise conditionally independent nonresponse (ICIN) mechanisms (Sadinle and Reiter, 2017), which are known to have identifiable parameters.
The imputation and estimation algorithm in Tang et al. (2024) does not use the design weights for the unit nonrespondents. Instead, they replace the design weights for unit nonrespondents with a constant that makes $\sum_{i=1}^{N}w_{i}=N$; that is, they base imputations and completed-data inferences on, for all units $i\in\mathcal{S},$
$w_{i}=\left\{\begin{array}[]{lr}w_{i}^{d}&\text{if }U_{i}=0\\ \frac{N-\sum_{k\in\mathcal{S}}w_{k}^{d}}{\sum_{k\in\mathcal{S}}U_{k}}&\text{if}U_{i}=1.\end{array}\right.$ (2.4)
For any unit $i\in\mathcal{S}$ and $j=1,\dots,p$, let $x_{ij}^{\star}=x_{ij}$ when $(R_{ij}=0,U_{i}=0)$, and let $x_{ij}^{\star}$ be an imputed value when $R_{ij}=1$ or $U_{i}=1$. For any $X_{j}\in\mathcal{A}$, Akande and Reiter (2022), Tang et al. (2024), and Yang and Reiter (2025) suppose the imputations to adhere to (2.1) – (2.3), with the additional constraint that
$\hat{T}_{jc}^{*}=\sum_{i\in\mathcal{S}}w_{i}I(x_{ij}^{\star}=c)\sim N(T_{jc},V_{jc}).$ (2.5)
Here, $V_{jc}$ is set by the analyst and determines how closely $\hat{T}_{jc}^{*}$ matches $T_{jc}$ in any completed dataset. As we also assume (2.5), we discuss setting $V_{jc}$ in Section 3.
3 Methodology
We now modify the hybrid missingness MD-AM model to use the design weights for all sampled individuals, including unit nonrespondents. For convenience, we presume that $X_{j}\in\mathcal{A}$ for $j=1,\dots,k<p$, and the remaining $p-k$ variables do not have auxiliary margins.
3.1 Variables with Item Nonresponse
We first create multiple imputations for all values of $x_{ij}$ missing due to item nonresponse for units with $U_{i}=0$. Following Yang and Reiter (2025), we implement multiple imputation by chained equations (MICE) using only units with $U_{i}=0$. The specification of the MICE algorithm does not require any extraordinary considerations. In our simulations, we use the “mice” package in R (Van Buuren, 2018), following its default settings including the ordering of variables. Each conditional model is a function of all other survey variables except $X_{j}$, including possibly interaction terms. The conditional models also could include subsets of $(R_{1},\dots,R_{j-1},R_{j+1},\dots,R_{p})$ as predictors; these are fully observed for units with $U_{i}=0$. We run the MICE procedure to create $L$ completed datasets for the units with $U_{i}=0$.
3.2 Variables with Margins for Unit Nonrespondents
After imputing values for item nonresponse, in each completed dataset we impute the unit nonrespondents’ values for all $X_{j}\in\mathcal{A}$. The analyst orders these variables based on ease of modeling. For notational convenience, we suppose that $X_{1}$ is imputed first, $X_{2}$ is imputed second, and so on sequentially until $X_{k}.$
To begin, for each unit $i$ with $U_{i}=1$ and in each of the $L$ completed datasets, the analyst sets an initial probability distribution for imputing $x_{i1}$. For example, in each completed dataset, the analyst estimates a (multinomial) logistic regression of $X_{1}$ on some function of $W$ and uses the predicted probabilities derived from the estimated model. Alternatively, the analyst could regress $X_{1}$ on an intercept only or compute marginal probabilities of $X_{1}$ using survey-weighted ratio estimators. For these two alternatives, the initial probabilities are identical for all unit nonrespondents. This approach may be especially appropriate when the analyst does not have strong reason to expect associations between the design weights and the missing values of $X_{1}$ for unit nonrespondents. Regardless of how it is derived, we refer to the initial distribution for imputing $x_{i1}$ as its working distribution, notated as $\{p_{i1c}:c=1,\dots,m_{1};U_{i}=1;i\in\mathcal{S}\}.$ Because of imputation for item nonresponse, the working distribution for any unit can differ across completed datasets; however, for convenience, we forego notation designating completed datasets. The analyst performs the computations of this section for each of the $L$ completed datasets.
Imputing each missing $x_{i1}$ using its working distribution does not utilize the information in $\mathcal{A}_{1}$. Thus, the resulting imputations could generate completed-data Horvitz and Thompson (1952) estimates that are far from the known totals for some levels of $X_{1}$. We therefore modify $\{p_{i1c}\}$ to ensure the resulting imputations result in reasonable completed-data point estimates of these totals.
To do so, we first simulate plausible values of $\hat{T}_{1c}$ for each level $c$ by appealing to large-sample central limit theorems. For $c=1,\dots,(m_{1}-1)$, we sample $\hat{T}_{1c}$ from $N(T_{1c},V_{1c})$ and set $\hat{T}_{1m_{1}}=N-\sum_{c=1}^{m_{1}-1}\hat{T}_{1c}$. We then find modified probabilities $\{\tilde{p}_{i1c}:c=1,\dots,m_{1};U_{i}=1;i\in\mathcal{S}\}$ so that, for $c=1,\dots,m_{1}$, the expectation of each completed-data estimator $\hat{T}_{1c}^{\star}$ from (2.5) over imputations for unit nonresponse approximately equals $\hat{T}_{1c}$. Put another way, recognizing that
$E\left(\sum_{i\in\mathcal{S}}w_{i}I(x_{ij}^{\star}=c)\right)$ $=$ $\sum_{i\in\mathcal{S}}w_{i}I(x_{ij}^{\star}=c)I(U_{i}=0)+E(\sum_{i\in\mathcal{S}}w_{i}I(x_{ij}^{\star}=c)I(U_{i}=1))$
$=$ $\sum_{i\in\mathcal{S}}w_{i}I(x_{ij}^{\star}=c)I(U_{i}=0)+\sum_{i\in\mathcal{S}}w_{i}\tilde{p}_{i1c}I(U_{i}=1),$
we want
$\sum_{i\in\mathcal{S}}w_{i}\tilde{p}_{i1c}I(U_{i}=1)$ $=$ $\hat{T}_{1c}-\sum_{i\in\mathcal{S}}w_{i}I(x_{i1}^{*}=c)I(U_{i}=0).$
To keep the imputation probabilities tied to the working distribution, for $c=1,\dots,m_{1}-1$, we set $\tilde{p}_{i1c}=f_{1c}p_{i1c}$, where each constant $f_{1c}>0$ is given by
$f_{1c}=\frac{\hat{T}_{1c}-\sum_{i\in\mathcal{S}}w_{i}I(x_{i1}^{*}=c,U_{i}=0)}{\sum_{i\in\mathcal{S}}p_{i1c}w_{i}I(U_{i}=1)}.$ (3.1)
We impute $x_{i1}^{*}$ for each unit with $U_{i}=1$ using a random draw from the adjusted probability mass function, $\{\tilde{p}_{i1c}:c=1,\dots,m_{1}\}$, where $\tilde{p}_{i1m_{1}}=1-\sum_{c=1}^{m_{1}-1}f_{1c}{p}_{i1c}$. If $\sum_{c=1}^{m_{1}-1}f_{1c}{p}_{i1c}>1$, we compute (3.1) for $c=1,\dots,m_{1}$ and set $\tilde{p}_{i1c}=f_{1c}{p}_{i1c}/\sum_{c=1}^{m_{1}}f_{1c}{p}_{i1c}$.
When each $p_{i1c}$ equals some common value $p_{1c}$ for all units with $U_{i}=1$, each $\tilde{p}_{i1c}=f_{1c}p_{1c}=\tilde{p}_{1c}$ is given conveniently by
$\tilde{p}_{1c}=\frac{\hat{T}_{1c}-\sum_{i\in\mathcal{S}}w_{i}I(x_{i1}^{*}=c,U_{i}=0)}{\sum_{i\in\mathcal{S}}w_{i}I(U_{i}=1)}.$ (3.2)
We now turn to imputing $X_{2}\in\mathcal{A}$ for unit nonrespondents. We follow a similar strategy: in each completed dataset, start by setting each unit nonrespondent’s working distribution for imputation of $x_{i2}^{*}$ given $x_{i1}^{*}$, then adjust the probabilities to make $(\hat{T}_{21}^{*},\dots,\hat{T}_{2m_{2}}^{*})$ approximately match a sampled value of $(\hat{T}_{21},\dots,\hat{T}_{2m_{2}}).$
Let $p_{i2c}=\Pr(X_{i2}=c|x_{i1}^{*}=d,w_{i})$ be the working probability that $X_{2}=c$ for unit $i$, given its imputed $x_{i1}^{*}=d$ and $w_{i}$. We determine these from a (multinomial) logistic regression of $X_{2}$ on $X_{1}$ and possibly some function of $W$, with coefficients estimated from the completed data for units with $U_{i}=0$. If we disregard $W$, $p_{i2c}=\Pr(X_{2}=c|X_{1}=d)$ for $c=1,\dots,m_{2}$ and $d=1,\dots,m_{1}$; that is, the conditional probability is the same for all units with a common value of $x_{i1}^{*}=d$.
For $c=1,\dots,m_{2}-1$, we sample a plausible value of $\hat{T}_{2c}$ by drawing from a $N(T_{2c},V_{2c})$, setting the estimated total for the final level as $\hat{T}_{2m_{2}}=N-\sum_{c=1}^{m_{2}-1}\hat{T}_{2c}$. Akin to (3.1), we then find suitable constants $f_{2c}>0$ to adjust each $p_{i2c}$. We have
$f_{2c}=\frac{\hat{T}_{2c}-\sum_{i\in\mathcal{S}}w_{i}I(x_{i2}^{*}=c,U_{i}=0)}{\sum_{i\in\mathcal{S}}p_{i2c}w_{i}I(U_{i}=1)}.$ (3.3)
We impute each $x_{i2}^{*}$ given $x_{i1}^{*}$ for units with $U_{i}=1$ using a random draw from the adjusted probability mass function $\{\tilde{p}_{i2c}:c=1,\dots,m_{2}\}$, where $\tilde{p}_{i2c}=f_{2c}{p}_{i2c}$ for $c=1,\dots,m_{2}-1$ and $\tilde{p}_{i2m_{2}}=1-\sum_{c=1}^{m_{2}-1}f_{2c}{p}_{i2c}$. If $\sum_{c=1}^{m_{2}-1}f_{2c}{p}_{i2c}>1$, we compute (3.3) for $c=1,\dots,m_{2}$ and set $\tilde{p}_{i2c}=f_{2c}{p}_{i2c}/\sum_{c=1}^{m_{2}}f_{2c}{p}_{i2c}$. This process ensures the completed-data Horvitz and Thompson (1952) estimates approximately match the sampled $(\hat{T}_{21},\dots,\hat{T}_{2m_{2}})$ in expectation, while also incorporating relationships between $X_{2}$ and $(X_{1},W)$ implied in the working probabilities. We independently repeat the estimation and imputation process in each completed dataset.
When $\{p_{i2c}\}$ derives from a logistic regression of $X_{2}$ on $X_{1}$ (and possibly $W$), using $\{\tilde{p}_{i2c}\}$ is equivalent to adjusting the intercept in that estimated regression, leaving other coefficients alone. Essentially, this implies a model where the log-odds for $X_{2}$ are the same for unit respondents and nonrespondents, per the hybrid missingness MD-AM assumption in (2.2) of no interaction between $U$ and $X_{2}$.
When disregarding $W$ in the logistic regression for $X_{2}$ on $X_{1}$—so that within any completed dataset, $p_{i2c}=\Pr(X_{2}=c|x_{1}=d,U=1)$ for all unit nonrespondents with $x_{i1}^{*}=d$ in that dataset—we can find $\{\tilde{p}_{2c}\}$ by solving a system of equations implied by (2.2). Specifically, we encode the constant log-odds assumption as a set of $(m_{1}-1)(m_{2}-1)$ equations. For $c=2,\dots,m_{2}$ and $d=2,\dots,m_{1}$, we have
$\text{log}\left(\frac{\Pr(X_{2}=c|X_{1}=d,U=1)}{\Pr(X_{2}=1|X_{1}=d,U=1)}\right)-\text{log}\left(\frac{\Pr(X_{2}=c|X_{1}=1,U=1)}{\Pr(X_{2}=1|X_{1}=1,U=1)}\right)$
$=\text{log}\left(\frac{\Pr(X_{2}=c|X_{1}=d,U=0)}{\Pr(X_{2}=1|X_{1}=d,U=0)}\right)-\text{log}\left(\frac{\Pr(X_{2}=c|X_{1}=1,U=0)}{\Pr(X_{2}=1|X_{1}=1,U=0)}\right).$ (3.4)
We can estimate the conditional probabilities for units with $U_{i}=0$ via survey-weighted ratio estimators using the completed data; for example,
$\Pr(X_{2}=c|X_{1}=d,U=0)=\frac{\sum_{i\in\textit{S}}w_{i}I(x_{i1}^{*}=d,x_{i2}^{*}=c,U_{i}=0)}{\sum_{i\in\textit{S}}w_{i}I(x_{i1}^{*}=d,U_{i}=0)}.$ (3.5)
We also encode the condition (2.5) on $(\hat{T}_{21}^{*},\dots,\hat{T}_{2m_{2}}^{*})$ as a set of $m_{2}$ equations. We have
$\sum_{d=1}^{m_{1}}\sum_{i\in\mathcal{S}}\Pr(X_{2}=c|X_{1}=d,U=1)w_{i}I(x_{i1}^{*}=d,x_{i2}^{*}=c,U_{i}=1)=\hat{T}_{2c}-\sum_{i\in\mathcal{S}}w_{i}I(x_{i2}^{*}=c,U_{i}=0).$ (3.6)
In total, we have $m_{1}(m_{2}-1)$ equations to solve for the $m_{1}(m_{2}-1)$ conditional probabilities, $\{\Pr(X_{2}=c|X_{1}=d,U=1):c=2,\dots,m_{2};d=1,\dots m_{1}\}$, which we use for imputation.
To illustrate, we present the computations for (3.4) and (3.6) for binary $X_{2}$ and $X_{1}$. We use these in the simulations of Section 4. First, we compute $\Pr(X_{2}=1|X_{1}=1,U=0)$ and $\Pr(X_{2}=1|X_{1}=0,U=0)$ using (3.5). To ease notation, define
$\hat{\beta}_{1,obs}$ $=$ $\text{logit}(\Pr(X_{2}=1|X_{1}=1,U=0))-\text{logit}(\Pr(X_{2}=1|X_{1}=0,U=0))$ (3.7)
$\hat{\beta}_{0,obs}$ $=$ $\text{logit}(\Pr(X_{2}=1|X_{1}=1,U=0))-\hat{\beta}_{1,obs}.$ (3.8)
Define the unobserved quantities $\hat{\beta}_{1,mis}$ and $\hat{\beta}_{0,mis}$ for the unit nonrespondents analogously, replacing $U_{i}=0$ with $U_{i}=1$ in (3.7) and (3.8). We set the log odds ratios equal for the unit respondents and nonrespondents, that is, we set $\hat{\beta}_{1,mis}=\hat{\beta}_{1,obs}.$ We also set
$\frac{\sum_{i\in\mathcal{S}}w_{i}I(x_{i1}^{*}=1,U_{i}=1)}{1+\exp(-(\hat{\beta}_{0,mis}+\hat{\beta}_{1,mis}))}+\frac{\sum_{i\in\mathcal{S}}w_{i}I(x_{i1}^{*}=0,U_{i}=1)}{1+\exp(-\hat{\beta}_{0,mis})}\\ =\hat{T}_{2}-\sum_{i\in\mathcal{S}}w_{i}I(x_{i2}^{*}=1,U_{i}=0).$
To determine $\hat{\beta}_{0,mis}$, we numerically solve the system of nonlinear equations using the modified Powell method as implemented in the R package “nleqslv” (Hasselman, 2018). The imputation probabilities are then
$\Pr(X_{2}=1|x_{i1}^{*}=1,U_{i}=1)$ $=$ $\frac{1}{1+\exp[-(\hat{\beta}_{0,mis}+\hat{\beta}_{1,mis})]}$ (3.9)
$\Pr(X_{2}=1|x_{i1}^{*}=0,U_{i}=1)$ $=$ $\frac{1}{1+\exp(-\hat{\beta}_{0,mis})}.$ (3.10)
When $k>2$, we can apply the process used in (3.3) to each $X_{j}\in\mathcal{A}$. For example, if $X_{3}\in\mathcal{A}$, the analyst specifies a set of working probabilities, $p_{i3c}=\Pr(X_{i3}=c|x_{i1}^{*},x_{i2}^{*},w_{i})$ for $c=1,\dots,m_{3}$ via a (multinomial) logistic regression of $X_{3}$ on $(X_{1},X_{2})$ and possibly $W$. The analyst samples values of $(\hat{T}_{31},\dots,\hat{T}_{3m_{3}})$ and finds the values of $f_{3c}$ using expressions analogous to (3.3), replacing quantities based on $X_{2}$ with those based on $X_{3}$. When the working probabilities are identical across nonresponding units, analysts can solve a system of equations akin to those in (3.4) and (3.6).
3.3 Variables Without Margins for Unit Nonrespondents
After creating $L$ partially completed datasets using Section 3.1 and 3.2, we impute values $x_{ij}^{*}$ for unit nonrespondents for all $X_{j}\notin\mathcal{A}$. We use the random hot deck imputation procedure developed by Yang and Reiter (2025). We briefly summarize the procedure here, closely following their presentation.
For any unit $i\in\mathcal{S}$ with $U_{i}=0$, in any completed dataset, let $\boldsymbol{x}_{i}^{\mathcal{A}}=\{x_{ij}:X_{j}\in\mathcal{A};j=1,\dots,p;U_{i}=0\}$ be the values of the (possibly imputed) survey variables in the completed data for those $X_{j}\in\mathcal{A}$. Similarly, for any unit $i^{\prime}\in\mathcal{S}$ with $U_{i^{\prime}}=1$, let $\boldsymbol{x}_{i^{\prime}}^{\mathcal{A}*}=\{x_{i^{\prime}j}^{*}:X_{j}\in\mathcal{A};j=1,\dots,p;U_{i}=1\}.$ For each unit $i^{\prime}$ with $U_{i^{\prime}}=1$, in each completed dataset we construct its donor set, $\mathcal{D}_{i^{\prime}}=\{(x_{i1},\dots,x_{ip}):\boldsymbol{x}_{i}^{\mathcal{A}}=\boldsymbol{x}_{i^{\prime}}^{\mathcal{A}*},U_{i}=0,i\in\mathcal{S}\}$. We randomly sample one record $i$ from $\mathcal{D}_{i^{\prime}}$ and append its $\{(x_{i1},\dots,x_{ik}):X_{j}\notin\mathcal{A}\}$ to $\boldsymbol{x}_{i^{\prime}}^{\mathcal{A}*}$ to make the full imputation $\boldsymbol{x}_{i^{\prime}}^{*}$ for unit $i^{\prime}$. We apply this procedure for all units in $\mathcal{S}$ with $U_{i^{\prime}}=1$, resulting in a completed dataset for all $n$ units. We repeat this process in each of the $L$ datasets.
4 Simulation Studies
To demonstrate the methodology, we mimic the simulation study in Yang and Reiter (2025). They found that, by leveraging auxiliary margins, multiple imputation via the hybrid missingness MD-AM model can offer more accurate inferences than multiple imputation that does not utilize $\mathcal{A}$. However, they use (2.4) in lieu of the design weights for unit nonrespondents, whereas we use them per Section 3.
We generate $\mathcal{P}$ comprising $N=3,373,378$ units. This is the number of individuals in the 2022 American Community Survey data file used by Yang and Reiter (2025). We make one of the variables in this file a size variable $Z=(z_{1},\dots,z_{N})$; this is observed for all $N$ individuals. To generate a $\mathcal{S}$, we sample records from $\mathcal{P}$ independently with $\pi_{i}=1/10z_{i}$, where $i=1,\dots,N$. This Poisson sampling design results in approximately $n=6000$ units in any $\mathcal{S}$. Let $W=10Z$ be the survey weights.
For each unit $i=1,\dots,N$, we generate four binary variables $(x_{i1},x_{i2},x_{i3},x_{i4})$ and two continuous variables $(x_{i5},x_{i6})$. The data generation model is displayed in (4.1)–(4.5).
$U$ $\sim\text{Bernoulli}(\pi_{U})\quad\text{logit}(\pi_{U})=\nu_{0}$ (4.1)
$X_{1}|U$ $\sim\text{Bernoulli}(\pi_{x_{1}})\quad\text{logit}(\pi_{x_{1}})=\omega_{10}+\omega_{11}W+\theta_{1}U$ (4.2)
$X_{2}|X_{1},U$ $\sim\text{Bernoulli}(\pi_{x_{2}})\quad\text{logit}(\pi_{x_{2}})=\omega_{20}+\omega_{21}X_{1}+\theta_{2}U$ (4.3)
$X_{j}|X_{1},\dots,X_{j-1},U$ $\sim\text{Bernoulli}(\pi_{x_{j}})\quad\text{logit}(\pi_{x_{j}})=\omega_{j0}+\sum_{k=1}^{j-1}\omega_{jk}X_{k}\,\,\,\textrm{for }j=3,4$ (4.4)
$X_{j}|X_{1},\dots,X_{j-1},U$ $\sim\text{Normal}(\omega_{j0}+\sum_{k=1}^{j-1}\omega_{jk}X_{k},\sigma_{j}^{2})\,\,\,\textrm{for }j=5,6.$ (4.5)
We set $\nu_{0}=-1.2$ for a unit nonresponse rate around 23%. We consider $\theta_{1}\in\{-2,-0.5\}$. The marginal distribution of $X_{1}$ for unit respondents and nonrespondents is substantially different when $\theta_{1}=-2$ and modestly different when $\theta_{1}=-0.5$. We use $\theta_{2}=-2$, which ensures the unit respondents and nonrespondents have different marginal distributions of $X_{2}$. The remaining parameter values are available in Appendix A.
We independently sample 500 datasets. For each sample $\mathcal{S}$, we generate the unit nonrespondents by sampling Bernoulli random draws of $U_{i}$, where $i\in\mathcal{S}$, from the distribution of $U|X_{1},X_{2}$ implied by (4.1)–(4.3). We make $\boldsymbol{x}_{i}$ completely missing for any unit in $\mathcal{S}$ where the drawn $U_{i}=1$. For item nonresponse, we use the ICIN mechanisms
$R_{j}|X_{1},X_{2},X_{3},X_{4},X_{5},X_{6},U=0\sim\text{Bernoulli}(\pi_{R_{j}}),\,\,\,\text{logit}(\pi_{R_{j}})=\phi_{j0}+\sum_{k\neq j}\phi_{jk}X_{k}$ (4.6)
for $X_{j}$ where $j\in\{2,3,4,6\}$, and we let $X_{1}$ and $X_{5}$ be fully observed. We blank every $x_{ij}$ for which the sampled $R_{ij}=1$. We set parameters in (4.6) so that approximately 25% of each eligible survey variable is missing for units with $U_{i}=0$. Specific parameter values are available in Appendix A.
We let $\mathcal{A}$ include $T_{1}=\sum_{i=1}^{N}x_{i1}$ and $T_{2}=\sum_{i=1}^{N}x_{i2}$, which we use for multiple imputation via the MD-AM models of Section 3. Using Section 3.1, we generate $L=10$ completed datasets for item nonresponse using a bespoke implementation of the “mice” package in R. Using Section 3.2, we consider two sets of working probabilities for $X_{1}$. The first allows $p_{i1c}$ to vary across units: each $p_{i1c}$ equals the unit’s predicted probability from the logistic regression of $X_{1}$ on $W$ estimated using the units with $U_{i}=0$. The second fixes $p_{i1c}=p_{1c}$ across units and uses (3.2). For both we use (3.9) and (3.10) to impute $X_{2}$. For the variances $V_{1}$ and $V_{2}$ used to sample $\hat{T}_{1}\sim N(T_{1},V_{1})$ and $\hat{T}_{2}\sim N(T_{2},V_{2})$, we use “mice” to make one completed dataset for all of $\mathcal{S}$ and compute the expressions for the unbiased variance estimators for $\hat{T}_{1}$ and $\hat{T}_{2}$ under Poisson sampling. Using Section 3.3, we create donor pools for the hot deck by matching on $(X_{1},X_{2})$.
After generating $L=10$ completed datasets, we use the inferential methods in Rubin (1987) for the estimands in Yang and Reiter (2025), including marginal totals for $(X_{1},\dots,X_{6})$ and several probabilities involving multiple categorical variables. In each completed dataset, we compute Horvitz and Thompson (1952) point and variance estimators for a Poisson sampling design based on the design weights for all sampled records. For comparisons, we also use the imputation strategy in Yang and Reiter (2025) that does not utilize $\mathcal{A}$. It uses a bespoke implementation of “mice” to impute missing items and randomly samples whole records with replacement as imputations for unit nonresponse. We refer to this method as IH.
Figure 4.1 summarizes the relative root mean squared errors (RMSEs) of the point estimates and empirical coverage rates of 95% multiple imputation confidence intervals for the three methods. Results in tabular form are at https://github.com/kevinxu47/MDAM. Both MD-AM models produce accurate point estimates, mitigating the bias from the nonignorable nonresponse evident in IH. The coverage rates for both MD-AM models tend to exceed 95% due to over-estimation of variances. As discussed by Yang and Reiter (2025), this is because the combining rules of Rubin (1987) do not account for using known totals in imputation. Overall, the results suggest that the MD-AM models enable analysts to incorporate known marginal totals in multiple imputation when using all sampled units’ design weights.
Figure 4.1: Simulated relative RMSEs and coverage rates for population totals, two-way conditional probabilities, and joint probabilities for MD-AM and IH models. Left panel compares IH to the MD-AM model with fixed working probabilities. Right panel displays the two MD-AM models. Top row uses $\theta_{1}=-2$. Bottom row uses $\theta_{1}=-0.5.$
Appendix A Simulation Parameters
The parameter values used in the simulations of Section 4 are identical to those in Yang and Reiter (2025). We set $\nu_{u}=-1.2$; $(\omega_{10},\omega_{11})=(0.06,-0.0002)$ and $\theta_{1}\in(-2,-0.5)$; $(\omega_{20},\omega_{21})=(0.2,0.4)$ and $\theta_{2}=-2$; $(\omega_{30},\omega_{31},\omega_{32})=(0.2,0.3,0.1)$; $(\omega_{40},\dots,\omega_{43})=(0.2,0.4,0.4,0.1)$; $(\omega_{50},\dots,\omega_{54})=(0.4,1.2,-0.9,0.1,0.2)$ and $\sigma_{5}=0.5$; $(\omega_{60},\dots,\omega_{65})=(0.4,1.2,-0.9,0.1,-0.1,0.1)$ and $\sigma_{6}=0.5$; $(\phi_{j0},\dots,\phi_{j5})=(-1.4,0.1,0.1,0.1,0.1,0)$ for $j\in\{2,3,4\}$,; and, $(\phi_{60},\dots,\phi_{65})$$=(-1.4,0.1,0.1,0.1,0.1,0.1)$.